Wellington Chocolate Company uses activity-based costing (ABC). The controller identified two activities and their budgeted costs: Setting up equipment $1,026,000 Other overheard $4,104,000 Setting up equipment is based on setup hours, and other overhead is based on oven hours. Wellington produces two products, Fudge and Cookies. Information on each product is as follows: Fudge Cookies Units produced 8,000 445,000 Setup hours 15,200 3,800 Oven hours 3,800 26,600 Required: Round your answers to the nearest whole dollar, unless otherwise directed.

1. Calculate the activity rate for (a) setting up equipment and (b) other overhead. a. Setting up equipment $ per setup hour b. Other overhead $ per oven hour

2. How much total overhead is assigned to Fudge using ABC? $

3. What is the unit overhead assigned to Fudge using ABC? Round to the nearest cent. $per unit

4. Now, ignoring the ABC results, calculate the plantwide overhead rate, based on oven hours. Round to the nearest cent. $ per oven hour

5. How much total overhead is assigned to Fudge using the plantwide overhead rate? $

6b. What is the difference in total overhead assigned to fudge under the two methods? $
Answer #1

Answer :-

1)The activity rate for (a) setting up equipment are as follows-

Activity rate for setting up equipment = Setting up equipment cost / Total setup hours

Given ,Settings up equipment cost = $1,026,000

Total setup hours = setup hours of Fudge + setup hours of Cookies

Total setup hours = 15,200 + 3,800 = 19,000

Activity rate for setting up equipment = $1,026,000 / 19,000

Activity rate for setting up equipment = $54 per setup hours

The activity rate for (b) other overhead are as follows :-

Activity rate for other overhead = Other overhead costs / Total oven hours

Given , Other overhead cost = $4,104,000

Total oven hours = Oven hours of Fudge + Oven hours of Cookies

Total oven hours = 3,800 + 26,600 = 30,400

Activity rate for other overhead = $4,104,000/ 30,400

Activity rate for other overhead = $135

2)Total overhead assigned to Fudge using ABC method is as follows :-

Total overhead = (Activity rate for setting up equipment × Fudge setup hours ) +( Activity rate for other overhead × Fudge Oven hours)

Total overhead = ($54 × 15,200) + ( $135 × 3800)

Total overhead= $1,333,800

Total overhead assigned to Fudge using ABC method is $1,333,800.

3)The unit overhead assigned to Fudge using ABC is :-

Unit overhead = Total overhead assigned to Fudge / No of unit Produced by Fudge

No.of unit produced by Fudge = 8000 units

Unit overhead = $1,333,800/ 8000

Unit overhead= $166,725 or $167

Unit overhead assigned to Fudge using ABC method is $167

4) The plantwide overhead rate, based on oven hours is as follows :-

Plantwide overhead rate = Total overhead costs / Total oven hours

Total overhead costs = Setting up equipment cost + Other overhead cost

Total overhead costs = $ 1,026,000 + $4,104,000

Total overhead costs = 5,130,000

Total oven hours = 30,400 (see Calculation above)

Plantwide overhead rate = 5,130,000/ 30,400

Plantwide overhead rate = $168.75

5) Total overhead is assigned to Fudge using the plantwide overhead rate :-

Total overhead of Fudge = Planttwide overhead rate × Oven hours of fedge

Total overhead of Fudge = $168.75 × 3,800

Total overhead of Fudge= $641,250

6)b) It above calculation we have find that their huge difference between the Total overhead of Fudge using ABC method and Plantwide overhead rate is because of setup hours. In ABC method we have consider the setup hour but Plantwide overhead rate method is based on oven hour.